# BAN: Detecting Backdoors Activated by Adversarial Neuron Noise

### v1

@@ Abstract
Backdoor attacks on deep learning represent a recent threat that has gained significant attention in the research community. Backdoor defenses are mainly based on backdoor inversion, which has been shown to be generic, model-agnostic, and applicable to practical threat scenarios. State-of-the-art backdoor inversion recovers a mask in the feature space to locate prominent backdoor features, where benign and backdoor features can be disentangled. However, it suffers from high computational overhead, and we also find that it overly relies on prominent backdoor features that are highly distinguishable from benign features. To tackle these shortcomings, this paper improves backdoor feature inversion for backdoor detection by incorporating extra neuron activation information. In particular, we adversarially increase the loss of backdoored models with respect to weights to activate the backdoor effect, based on which we can easily differentiate backdoored and clean models. Experimental results demonstrate our defense, BAN, is 1.37$\times$ (on CIFAR-10) and 5.11$\times$ (on ImageNet200) more efficient with 9.99% higher detect success rate than the state-of-the-art defense BTI-DBF [43]. Our code and trained models are publicly available. [^1] https://anonymous.4open.science/r/ban-4B32
@@ 1 Introduction
  • We propose detecting Backdoors activated by Adversarial neuron Noise (BAN) to mitigate this generalization shortcoming by introducing neuron noise into feature space trigger inversion. BAN includes an adversarial learning process to incorporate neuron activation information into the inversion-based backdoor detection. Experimental results demonstrate that BAN is 1.37$\times$ (on CIFAR-10) and 5.11$\times$ (on ImageNet200) more efficient with 9.99% higher detect success rate than the state-of-the-art defense BTI-DBF [43].
  • We also exploit the neuron noise to further design a simple yet effective defense for removing the backdoor, such that we build a complete defense framework.
@@ 3.1 The Pipeline of Training Backdoor Models
For brevity, consider an $L$-layer fully connected network $f$ (similar principles apply to convolutional networks) that has $l=l_{1}+l_{2}+\cdots+l_{L}$ neurons. $\mathbf{x}_{n}\in\mathbb{R}^{d_{x}}$ and $y_{n}\in\{0,1\}^{d_{y}}$ are the $n_{\text{th}}$ image and its label in $d_{x}$ and $d_{y}$ dimensional spaces, respectively. The attacker creates a poisoned dataset $\mathcal{D}_{p}$ by poisoning generators $G_{X}$ and $G_{Y}$ for a subset of the whole training dataset, i.e., $\mathcal{D}_{p}=\mathcal{D}_{c}\cup\mathcal{D}_{b}$. $\mathcal{D}_{c}$ is the original clean data. $\mathcal{D}_{b}$ is the poisoned backdoor data, $\mathcal{D}_{b}=\{(\mathbf{x}^{\prime},y^{\prime})|\mathbf{x}^{\prime}=G_{X}(\mathbf{x}),y^{\prime}=G_{Y}(y),(\mathbf{x},y)\in\mathcal{D}-\mathcal{D}_{c}\}$. In all-to-one attacks, $G_{Y}(y)=y_{t}$, $y_{t}$ is the attacker-specified target class. In our experiments, we consider the dirty-label attack. In all-to-all attacks, ususally $G_{Y}(y)=(y+1)$ [43, 25, 39], which is also what we chose in our experiemnts. In the training stage, the backdoor is injected into the model by training with $\mathcal{D}_{p}$, i.e., minimizing the training loss on $\mathcal{D}_{p}$ to find the optimal weights and bias $(\mathbf{w}^{*},\mathbf{b}^{*})$:
@@ 3.3 Detection with Neuron Noise
Neuron Noise. In backdoor models, there are two types of neurons: benign and backdoor [20]. The backdoor neurons build a shortcut from $G_{X}(\mathbf{x})$ to $G_{Y}(y)$. Neuron noise is generated noises on neurons to maximize model loss on clean samples in an adversarial manner [22]. The noise on benign neurons evenly misleads prediction to all classes, while the noise on backdoor neurons tends to mislead prediction to the target label due to the backdoor shortcut, as shown in Figure 1. Therefore, backdoor models with noise behave differently from benign models with noise, as there are no backdoor neurons in the benign models.
@@
where $\odot$ is the element-wise product. This optimization divides the latent features into two parts by the mask while maintaining a relatively small size of $\mathbf{m}$. Note that the regularizer for $\mathbf{m}$ in Eq. (5) is necessary. Otherwise, $\mathbf{m}$ will become a dense matrix full of ones to focus on the positive part of Eq. (5) because maintaining only the positive part (without penalty of $|\mathbf{m}|$) already satisfies the optimization objective. Finally, we apply the negative mask $(1-\mathbf{m})$ on top of latent feature of $f$ to enhance the backdoor effect. The final output is:
@@ 4.1 The Performance of Backdoor Detection
Main Results. In Table 1, BAN shows better results on CIFAR-10 than all baselines, especially on advanced attacks. Results on other datasets are presented in Appendix C. We note that the advanced detection methods (FeatureRE, Unicorn, and BTI-DBF) perform worse than NC on simple attacks (BadNets and Blend). We hypothesize this is because the backdoor features generated by BadNets are less obvious on feature channels than advanced attacks in the feature space, such as WaNet and IAD. Figure 3 shows that BadNets features are weaker than features from advanced attacks using the feature mask in Eq. (5). Specifically, we optimize the feature mask to disentangle the benign and backdoor features for four attacks.
Figure 3: BadNet features are weaker in the feature space when we use the mask that disentangles the benign and backdoor features. Defenses that are biased towards large differences may not work in cases like BadNets.
Then, we compute two cross-entropy loss values using the positive mask ($\mathbf{m}$) and the negative mask ($1-\mathbf{m}$) for benign and backdoor features, respectively. The average loss values and standard deviation for four models for each attack are plotted in Figure 3. The negative loss value of BadNets is much smaller than others, which means BadNets features are weaker than others with regard to misleading the model to the backdoor target. Recent defenses usually add more regularizers to their losses and optimization objectives to counteract powerful backdoor attacks. These regularizers encourage the capturing of strong features but omit weak ones. Thus, recent advanced detection methods can perform worse on BadNets than NC.
Table 1: The detection results under different model architectures on CIFAR-10. The “Bd.” refers to the number of models that the defense identifies as backdoored. The “Acc.” refers to detection success accuracy. The best results are marked in bold. BTI-DBF* refers to an improved version (details in Section 3.4).
@@ 4.3 Defense against All-To-All Attacks
Previous works [37, 38] usually only consider all-to-one attacks, which limits the application in practical situations. In this section, we evaluate our fine-tuning method under three all-to-all attacks: WaNet-All [25], IAD-All [24], and Bpp-All [39]. FeatureRE is designed for all-to-one attacks, so we use target label 0 here for FeatureRE. BAN is capable of handling all-to-all attacks because we directly explore the neurons themselves instead of optimizing for the potential targeted label. Table 5 demonstrates the effectiveness of our method.
@@ 4.4 Evaluation under Adaptive Attack
Attackers may design a specific attack for defense when they know its details [33]. In this section, we evaluate BAN against two attacks that attempt to bypass the difference between backdoor and benign features: Adap-Blend [27] and SSDT attack [23]. Both Adap-Blend [27] and SSDT attack [23] try to obscure the difference between benign and backdoor latent features. Adap-Blend [27] achieves it by randomly applying 50% of the trigger to poison the training data, while SSDT attack [23] utilizes source-specific and dynamic-triggers to reduce the impact of triggers on samples from non-victim classes. The source-specific attack refers to backdoor triggers that mislead the model to the target class only when they are applied to victim class samples. Table 6 demonstrates our approach is resistant to these two attacks, while other methods fail. The reason is that the backdoor features of SSDT are close to benign features in the feature space. It is difficult for other methods to distinguish between backdoor and benign features created by SSDT. Our detection method directly analyzes the model itself using neuron noise, which captures the difference between backdoor and benign models concerning parameters. See Appendix C.3 for the results of our method against these adaptive attacks.
@@ B.4 BAN Settings
Detection. We use epsilon ($\epsilon$) to limit the noise added to neurons. Otherwise, the model weights are destroyed by the noise. The $\epsilon$ values are 0.3, 0.3, 0.2, and 0.1 for CIFAR-10, GTSRB, Tiny-ImageNet, and ImageNet200, respectively. $\epsilon$ values are decided according to the size of the images. We use smaller $\epsilon$ for larger datasets. We use the 30-step PGD algorithm to solve the optimization in Eq. (4) to find the noise, i.e., $\bm{\delta}$ and $\bm{\xi}$. We use SGD and the learning rate of $\epsilon/30$ for the PGD optimization. Then, we use Adam and the learning rate of 0.01 to search for 20 epochs for the feature mask using Eq. (5). The $\lambda_{1}$ for Eq. (5) equals 0.75. The two optimizations above use 1% of the training set. Note that this 1% of data is not used to train the backdoor models. The elements in the mask are clamped into continuous values between 0 and 1.
@@ C.4 Comparsion with ANP
Table 11: Comparsion with ANP [40] on CIFAR-10 using ResNet18.

### v2

@@ Abstract
Backdoor attacks on deep learning represent a recent threat that has gained significant attention in the research community. Backdoor defenses are mainly based on backdoor inversion, which has been shown to be generic, model-agnostic, and applicable to practical threat scenarios. State-of-the-art backdoor inversion recovers a mask in the feature space to locate prominent backdoor features, where benign and backdoor features can be disentangled. However, it suffers from high computational overhead, and we also find that it overly relies on prominent backdoor features that are highly distinguishable from benign features. To tackle these shortcomings, this paper improves backdoor feature inversion for backdoor detection by incorporating extra neuron activation information. In particular, we adversarially increase the loss of backdoored models with respect to weights to activate the backdoor effect, based on which we can easily differentiate backdoored and clean models. Experimental results demonstrate our defense, BAN, is 1.37$\times$ (on CIFAR-10) and 5.11$\times$ (on ImageNet200) more efficient with an average 9.99% higher detect success rate than the state-of-the-art defense BTI-DBF. Our code and trained models are publicly available at https://github.com/xiaoyunxxy/ban. Corresponding author.
@@ 1 Introduction
  • We propose detecting Backdoors activated by Adversarial neuron Noise (BAN) to mitigate this generalization shortcoming by introducing neuron noise into feature space trigger inversion. BAN includes an adversarial learning process to incorporate neuron activation information into the inversion-based backdoor detection. Experimental results demonstrate that BAN is 1.37$\times$ (on CIFAR-10) and 5.11$\times$ (on ImageNet200) more efficient with a 9.99% higher detect success rate than the state-of-the-art defense BTI-DBF [44].
  • We also exploit the neuron noise to further design a simple yet effective defense for removing the backdoor, such that we build a workable framework.
@@ 3.1 The Pipeline of Training Backdoor Models
For brevity, consider an $L$-layer fully connected network $f$ (similar principles apply to convolutional networks) that has $l=l_{1}+l_{2}+\cdots+l_{L}$ neurons. $\mathbf{x}_{n}\in\mathbb{R}^{d_{x}}$ and $y_{n}\in\{0,1\}^{d_{y}}$ are the $n_{\text{th}}$ image and its label in $d_{x}$ and $d_{y}$ dimensional spaces, respectively. The attacker creates a poisoned dataset $\mathcal{D}_{p}$ by poisoning generators $G_{X}$ and $G_{Y}$ for a subset of the whole training dataset, i.e., $\mathcal{D}_{p}=\mathcal{D}_{c}\cup\mathcal{D}_{b}$. $\mathcal{D}_{c}$ is the original clean data. $\mathcal{D}_{b}$ is the poisoned backdoor data, $\mathcal{D}_{b}=\{(\mathbf{x}^{\prime},y^{\prime})|\mathbf{x}^{\prime}=G_{X}(\mathbf{x}),y^{\prime}=G_{Y}(y),(\mathbf{x},y)\in\mathcal{D}-\mathcal{D}_{c}\}$. In all-to-one attacks, $G_{Y}(y)=y_{t}$, $y_{t}$ is the attacker-specified target class. In our experiments, we consider the dirty-label attack. In all-to-all attacks, ususally $G_{Y}(y)=(y+1)$ [44, 26, 40], which is also what we chose in our experiments. In the training stage, the backdoor is injected into the model by training with $\mathcal{D}_{p}$, i.e., minimizing the training loss on $\mathcal{D}_{p}$ to find the optimal weights and bias $(\mathbf{w}^{*},\mathbf{b}^{*})$:
@@ 3.3 Detection with Neuron Noise
Neuron Noise. In backdoor models, there are two types of neurons: benign and backdoor [21]. The backdoor neurons build a shortcut from $G_{X}(\mathbf{x})$ to $G_{Y}(y)$. Neuron noise is generated noises added on neurons to maximize model loss on clean samples in an adversarial manner [23]. The noise on benign neurons evenly misleads prediction to all classes, while the noise on backdoor neurons tends to mislead prediction to the target label due to the backdoor shortcut, as shown in Figure 1. Therefore, backdoor models with noise behave differently from benign models with noise, as there are no backdoor neurons in the benign models.
@@
where $\odot$ is the element-wise product. This optimization divides the latent features into two parts through the mask while maintaining a relatively small size of $\mathbf{m}$. Note that the regularizer for $\mathbf{m}$ in Eq. (5) is necessary. Otherwise, $\mathbf{m}$ will become a dense matrix full of ones to focus on the positive part of Eq. (5) because maintaining only the positive part (without penalty of $|\mathbf{m}|$) already satisfies the optimization objective. Finally, we apply the negative mask $(1-\mathbf{m})$ on top of latent feature of $f$ to enhance the backdoor effect. The final output is:
@@ 4 Experimental Results
Figure 3: BadNets features are weaker when using the mask to disentangle the benign and backdoor features. Defenses that are biased towards large differences may not work in cases like BadNets.
@@ 4.1 The Performance of Backdoor Detection
Main Results. In Table 1, BAN shows better results on CIFAR-10 than all baselines, especially on advanced attacks. Results on other datasets are presented in Appendix C. We note that the advanced detection methods (FeatureRE, Unicorn, and BTI-DBF) perform worse than NC on simple attacks (BadNets and Blend). We hypothesize this is because the backdoor features generated by BadNets are less obvious on feature channels than advanced attacks in the feature space, such as WaNet and IAD. Figure 3 shows that BadNets features are weaker than features from advanced attacks using the feature mask in Eq. (5). Specifically, we optimize the feature mask to disentangle the benign and backdoor features for four attacks. Then, we compute two cross-entropy loss values using the positive mask ($\mathbf{m}$) and the negative mask ($1-\mathbf{m}$) for benign and backdoor features, respectively. The average loss values and standard deviation for four models for each attack are plotted in Figure 3. The negative loss value of BadNets is much smaller than others, which means BadNets features are weaker than others with regard to misleading the model to the backdoor target. Recent defenses usually add more regularizers to their losses and optimization objectives to counteract powerful backdoor attacks. These regularizers encourage the capturing of strong features but omit weak ones. Thus, recent advanced detections can perform worse on BadNets than NC.
Table 1: The detection results under different model architectures on CIFAR-10. The “Bd.” refers to the number of models the defense identifies as backdoored. The “Acc.” refers to detection success accuracy. The best results are marked in bold. BTI-DBF* refers to an improved version (details in Section 3.4).
@@ 4.3 Defense against All-To-All Attacks
Previous works [38, 39] usually only consider all-to-one attacks, which limits the application in practical situations. In this section, we evaluate our fine-tuning method under three all-to-all attacks: WaNet-All [26], IAD-All [25], and Bpp-All [40]. FeatureRE is designed for all-to-one attacks, so we use target label 0 here for FeatureRE. FeatureRE is included to show that the all-to-one defense does not work in the all-to-all setting. BAN is capable of handling all-to-all attacks because we directly explore the neurons themselves instead of optimizing for the potential targeted label. Table 5 demonstrates the effectiveness of our method.
@@ 4.4 Evaluation under Adaptive Attack
Attackers may design a specific attack for defense when they know its details [34]. In this section, we evaluate BAN against two attacks that attempt to bypass the difference between backdoor and benign features: Adap-Blend [28] and SSDT attack [24]. Both Adap-Blend [28] and SSDT attack [24] try to obscure the difference between benign and backdoor latent features. Adap-Blend [28] achieves it by randomly applying 50% of the trigger to poison the training data, while SSDT attack [24] utilizes source-specific and dynamic-triggers to reduce the impact of triggers on samples from non-victim classes. The source-specific attack refers to backdoor triggers that mislead the model to the target class only when applied to victim class samples. Table 6 demonstrates our approach is resistant to these two attacks, while other methods fail. The reason is that the backdoor features of SSDT are close to benign features in the feature space. It is difficult for other methods to distinguish between backdoor and benign features created by SSDT. Our detection method directly analyzes the model itself using neuron noise, which captures the difference between backdoor and benign models concerning parameters. See Appendix C.3 for the mitigation results of our method against these adaptive attacks.
@@
4.5 Analysis on Prominent Features
We provide additional analysis of the phenomenon that backdoor features are more prominent for advanced attacks (WaNet, IAD, and Bpp) than weaker attacks (BadNets, Blend). Table 7 demonstrates that previous decoupling methods cannot easily pick up backdoor features from weak attacks, such as BadNets. In particular, when detecting without the $L_{1}$ regularizer (i.e., w/o norm), the negative feature loss of BadNets is high with a very large $L_{1}$ mask norm, while the Bpp has an even higher negative loss with a much smaller mask norm. The high negative loss of BadNets is actually from the sparse feature mask rather than backdoor features, i.e., there are too many zeros in (1 - $\mathbf{m}$). This indicates that BadNet backdoor features are less prominent than Bpp features, making it more challenging to decouple BadNets features.
Table 7: Loss values when feeding the benign or backdoor features into the final classification layer. The mask is optimized using Eq. (5) (w/ norm) or Eq. (7) (w/o norm). The shape of the feature mask is $512\times 4\times 4$, so the maximal $L_{1}$ norm is 8192.
Attack BA ASR $L_{1}$ mask norm Positive feature loss Negative feature loss
BadNets (w/ norm) 93.47 99.71 2258.90 0.21 0.26
BadNets (w/o norm) 93.47 99.71 8054.45 0.14 2.17
Blend (w/ norm) 94.60 100.00 2084.62 0.15 0.20
Blend (w/o norm) 94.60 100.00 8117.90 0.04 2.22
WaNet (w/ norm) 93.88 99.63 7400.97 0.06 2.34
WaNet (w/o norm) 93.88 99.63 7702.56 0.05 2.39
IAD (w/ norm) 93.82 99.64 7898.91 0.03 2.25
IAD (w/o norm) 93.82 99.64 7895.16 0.03 2.25
Bpp (w/ norm) 94.56 99.97 7147.68 0.09 2.80
Bpp (w/o norm) 94.56 99.97 7260.31 0.09 2.78
@@ 5 Conclusions and Future Work
Acknowledgments and Disclosure of Funding
This work was partly funded by the Dutch Research Council (NWO) through the PROACT project (NWA.1215.18.014) and the CiCS project of the research programme Gravitation (024.006.037).
@@ B.4 BAN Settings
Detection. We use epsilon ($\epsilon$) to limit the noise added to neurons. Otherwise, the model weights are destroyed by the noise. The $\epsilon$ values are 0.3, 0.3, 0.2, and 0.1 for CIFAR-10, GTSRB, Tiny-ImageNet, and ImageNet200, respectively. $\epsilon$ values are decided according to the size of the images. We use smaller $\epsilon$ for larger datasets. We use the 30-step PGD algorithm to solve the optimization in Eq. (4) to find the noise, i.e., $\bm{\delta}$ and $\bm{\xi}$. We use SGD and the learning rate of $\epsilon/30$ for the PGD optimization. Then, we use Adam and the learning rate of 0.01 to search for 20 epochs for the feature mask using Eq. (5). The $\lambda_{1}$ for Eq. (5) equals 0.75. The two optimizations above use 1% of the training set. Note that this 1% of data is not used to train the backdoor models. The elements in the mask are clamped into continuous values between 0 and 1. In Figure 5, we present an illustrative diagram of BAN.
@@ C.4 Comparsion with ANP
Table 12: Comparison with ANP [41] on CIFAR-10 using ResNet18.
@@
C.5 Different Poisoning Rates
BAN is effective against BadNets with different poisoning rates, as shown in Table 13. We use BadNets because it is relatively weak at a low poisoning rate (i.e., hard to be mitigated), while more advanced attacks may still be strong at a low poisoning rate.
Table 13: The performance of BAN fine-tuning under different poisoning rates of BadNets using CIFAR-10.
Poisoning Rate BA ASR pos. feature loss neg. feature loss Mitigated BA Mitigated ASR
0.01 93.48 98.69 0.38 0.17 92.07 2.73
0.05 93.37 99.41 0.37 0.35 92.06 1.97
0.10 90.98 100.00 0.35 2.06 90.29 2.17
0.15 90.32 100.00 0.39 2.23 90.16 1.71
0.20 89.34 100.00 0.44 2.43 90.39 1.01
0.25 88.09 100.00 0.56 2.81 89.55 1.54
0.30 86.09 100.00 0.62 3.13 88.83 1.08
0.40 82.39 100.00 0.67 3.51 88.75 1.67
0.50 77.83 99.97 0.84 4.27 86.87 3.56
C.6 Against Backdoors on the MLP
We evaluate the defense performance on simple model architecture. In particular, a 4-layer MLP is trained with benign samples and with BadNets. Table 14 demonstrates that BAN is effective on MLP, where the “num. to target” refers to the number of samples (in 5000 validation samples) that are classified as the backdoor target after our detection. We also find that the positive feature loss (i.e., benign feature loss) is very close to the negative loss (i.e., potential backdoor feature loss), which indicates that the backdoor features are more challenging to decouple from benign ones.
Table 14: The performance of BAN on the 4-layer MLP.
MLP BA ASR Pos. feature loss Neg. feature loss Num. to target Mitigated BA Mtigated ASR
Benign 53.24 - 1.95 2.29 419 - -
BadNets 47.04 100.00 2.03 2.34 3607 45.13 7.11
C.7 Discussion about Relationship between the Neuron Noise and Lipschitz Continuity
A small trigger that changes the output of a benign input into a malicious target label can be related to the high Lipschitz constant [45] and a neural network with high robustness tends to have a lower local Lipschitz constant. Moreover, a larger local Lipschitz constant implies steeper output around trigger-inserted points, leading to a smaller trigger effective radius making trigger inversion less effective [49]. Thus, the concepts of adversarial activation noise and Lipschitz continuity are related, and the local Lipschitz constant can serve as an upper bound for the trigger’s effective influence.
In addition, introducing theoretical tools like the Lipschitz constant for a backdoor defense may also be very tricky in practice because it needs approximation for implementation. For example, [45] evaluates the channel-wise Lipschitz constant by its upper bound but does not thoroughly discuss the relationship between the channel-wise Lipschitz constant and the network-wise Lipschitz constant, where the theorem of Lipschitz continuity really relies on. Recent work also mentions that empirically estimating the Lipschitz constant is hard from observed data, which usually leads to overestimation [5]. Methods relying on Lipschitz continuity may not require heavy computational load and are also related to our approach. Our method emphasizes more on fine-tuning with the guidance of neuron noise, rather than tuning the trained model.
C.8 Choosing the $\lambda_{1}$
As discussed in the method section, the mask norm ($L_{1}$ norm) with lambda in Eq. (5) is to ensure that the optimization objective is decoupling between benign and backdoor features. Without the constraint of the mask norm in Eq. (5), the optimization objective will be simply increasing the mask norm unless there are extremely strong backdoor features. The $\lambda_{1}$ value selection is implemented by checking the value of the mask norm. Table 15 provides an example of $\lambda_{1}$ selection on CIFAR-10, where the maximal mask norm is 8192. It can be observed that the mask is almost full of ones when $\lambda_{1}$ is smaller than 0.7, and the negative feature loss (backdoor feature) is high, based on which we can pick a value of $\lambda_{1}$ greater than 0.7. Note that the selection is unaware of potential backdoors.
Table 15: Feature loss values with different $\lambda_{1}$
$\lambda_{1}$ Mask norm Pos. feature loss Neg. feature loss
0.0 8188.62 0.27 2.30
0.1 8188.75 0.28 2.30
0.2 8184.30 0.27 2.30
0.3 8175.50 0.29 2.30
0.4 8152.40 0.25 2.30
0.5 8131.26 0.27 2.29
0.6 8055.07 0.21 2.27
0.7 7898.25 0.26 2.24
0.8 596.85 0.99 0.23
0.9 22.08 2.33 0.28
C.9 Swin Transformer
Table 16 provides experimental results of a 12-layer Swin transformer on CIFAR-10. For BadNets and Blend, we train the backdoored network using Adam as the optimizer. For WaNet, IAD, and Bpp, we use the default setting. All backdoored models can be detected by BAN.
Table 16: Results on Swin Transformer.
Attack No defense FT BTI-DBF(U) BAN Fine-tune
BA ASR BA ASR BA ASR BA ASR
BadNets 86.7 97.43 82.23 44.07 85.21 99.99 83.92 2.91
Blend 85.46 100.00 80.13 100.00 83.9 100.00 79.18 26.20
WaNet 78.25 31.03 75.63 2.3 73.83 2.61 75.28 3.1
IAD 76.35 88.71 74.58 54.63 74.29 61.16 76.26 9.8
Bpp 79.19 63.74 75.35 12.99 74.66 11.04 74.25 6.91
C.10 Clean Label Attack
We provide a clean-label backdoor experiment following [35], where we take default hyperparameters. Table 17 demonstrates the detection performance of 20 models, where we randomly select one model from the 20 for the mitigation experiment. Experimental results demonstrate that BAN is also effective against the clean-label attack.
Table 17: Clean label attack on CIFAR-10.
Attack BA ASR Detection Success Accuracy Mitigated BA Mitigated ASR
LC [35] 92.72 100.00 90.00 89.27 6.16
Table 18: Results on different target classes on CIFAR-10.
Target No defense BAN Fine-tune
BA ASR BA ASR
0 93.41 100.00 92.57 1.56
1 93.51 100.00 92.53 0.84
2 93.54 99.99 91.49 1.08
3 93.59 99.99 92.14 1.93
4 93.84 99.98 92.13 1.49
5 93.52 100.00 91.84 3.29
6 93.46 100.00 92.48 0.93
7 93.56 100.00 92.36 0.73
8 93.57 99.89 92.31 0.58
9 93.36 100.00 91.65 2.40
C.11 Semantically Similar Classes
We provide an additional analysis on image form semantically similar classes. From ImageNet200, we select class n02096294 (Australian terrier) as the target class, since there are 19 kinds of terriers in our ImageNet200 datasets, such as n02094433 (Yorkshire terrier) and n02095889 (Sealyham terrier). We train 10 backdoor networks using BadNets with different target classes. Table 18 demonstrates that BAN is effective against the backdoor regardless of semantically similar classes.